OpenStudy (anonymous):

Can you help me @mathstudent55 ? Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. A. has an x-intercept of 6 and a y-intercept of 3

OpenStudy (mathstudent55):

Slope-intercept form is \(y = mx + b\) Point-slope is \(y - y_1 = m(x - x_1) \) Standard form is \(ax + by = c\) Since you are given the y-intercept, it's easy to use the slope-intercept form. All we need to find is the slope. slope = \(m = \dfrac{y_2 - y_1}{x_2 - x_1} \) Start by using the two intercepts to find the slope. The x-intercept is (6, 0), and the y-intercept is (0, 3). Use those two points to find m, the slope. Then use b = 3 and the slope in the slope-intercept form of the equation.

OpenStudy (anonymous):

ok, thanks!!! :)

OpenStudy (mathstudent55):

wlcm